Circular No. 8060

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003V Further to IAUC 8053, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of an apparent supernova (mag 18.2) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Jan. 16.16, 16.18, and 16.20 UT, with confirming images from Jan. 25 showing the new object at mag 18.4. SN 2003V is located at α = 3^{h}05^{m}43^{s}.05, δ = −1^{o}24′12′′.8 (equinox 2000.0). This is 3′′ west and 2′′ north of the center of the apparent host galaxy. No point source is visible at this location in previous Palomar NEAT images taken on 2002 Dec. 7, 10, 12, and 13 (subtracted against earlier NEAT images from 2001 and 2002; limiting mag of 22.2 at S/N of 3).

FLARING OBJECT IN ORION CLUSTER K. Nakanishi and M. Saito, Nobeyama Radio Observatory, National

Astronomical Observatory of Japan; R. S. Furuya, Istituto Nazionale di Astrofisica, Osservatorio Astrofisico di Arcetri; H. Shinnaga, Harvard-Smith-sonian Center for Astrophysics; and M. Momose, Ibaraki University, report on follow-up observations of the radio flaring object in Orion cluster, using the Nobeyama Millimeter Array (NMA) at Nobeyama Radio Observatory: The object is located at α = 5^{h}35^{m}11^{s}.79, δ = −5^{o}21′49′′.2 (equinox 2000.0; estimated uncertainty 0′′.2), which agrees well with that reported on IAUC 8055. The NMA data give continuum flux densities of 39.1 ± 5.1 mJy at 146.969 GHz on Jan. 25.55 UT (resolution 1′′.3 × 0′′.9) and 22.4 ± 5.4 mJy at 97.782 GHz on Jan. 25.62 (resolution 2′′.5 × 1′′.2). We suggest that the object is/are (a) massive (proto)star(s) that might be involved in mass-ejection event(s). Further interferometric observations at millimeter wavelengths are planned, and observations at other wavelengths are urgently requested.”

SUPERNOVAE 2003P AND 2003T R. J. Foley, J. Graham, M. Ganeshalingam, and A. V. Filippenko,

University of California, Berkeley, report that inspection of CCD spectra (range 310–1000 nm), obtained on Jan. 28 UT with the Shane 3-m telescope at Lick Observatory, reveals that SN 2003P (IAUC 8057) is of type Ia, roughly 1 month past maximum brightness. SN 2003T (IAUC 8058) is of type II, perhaps a week after the explosion; hydrogen Balmer lines with P-Cyg profiles are superposed on a blue continuum, but the emission component of Hα greatly dominates over the absorption component.

2003 January 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green